OpenStudy (anonymous):

Can someone please help me in Algebra II? I don't understand the substitution method. I have a couple problems I need help with. Thanks!

OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y – z = 2 3x + y – 2z = –15 3x + 4y – 3z = –2

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@phi

OpenStudy (phi):

for more than 2 variables, I don't like substitution.... do you have to do substitution ?

OpenStudy (anonymous):

Yes, unfortunately...

OpenStudy (phi):

ok, can you solve for x in the first equation ?

OpenStudy (anonymous):

I don't know how to...

OpenStudy (anonymous):

There are two other variables... How would I solve for x?

OpenStudy (phi):

"move" the other terms to the right-hand side of the = sign to move the 2y add -2y to both sides to move the -z add +z to both sides what do you get ?

OpenStudy (anonymous):

2-2y+z

OpenStudy (phi):

yes, but you should write it as x = 2-2y+z because solving this mess means being very careful, and writing everything down (I lose track if I don't) now replace x with (2-2y+z) in the other two equations can you do that ?

OpenStudy (anonymous):

I think so

OpenStudy (amistre64):

or just multiply that by 3 an replace the whole 3x shebang

OpenStudy (anonymous):

3(2-2y+z)+y-2z=-15 and 3(2-2y+z)+4y-3z=-2

OpenStudy (phi):

next, for both equations , distribute the 3 and then combine "like terms" can you do that ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-5y+z=-9 and -2y=4

OpenStudy (anonymous):

-2y=-8?

OpenStudy (phi):

almost. I think after distributing the 3 you get (in the first equation) 6 -6y + 3z + y -2z = -15 combine -6y + y to get -5y combine 3z -2z to get +z 6 -5y + z = -15 add -6 to both sides to get -5y + z = -15-6 = -21 can you redo the 2nd equation ?

OpenStudy (phi):

yes, so you have -5y + z = -21 -2y = -8 now solve for y. I would use the 2nd equation (for obvious reasons)

OpenStudy (anonymous):

y=4

OpenStudy (phi):

now use y=4 in the other equation -5y + z = -21 and solve for z

OpenStudy (anonymous):

-1?

OpenStudy (phi):

finally use y=4 and z=-1 in x = 2-2y+z

OpenStudy (anonymous):

-3?

OpenStudy (phi):

how ?

OpenStudy (anonymous):

wait... is it 7?

OpenStudy (anonymous):

or -5..? I'm confused...

OpenStudy (phi):

don't lose the bubble... start with x= 2 - 2*y + z replace y with 4 x = 2 - 2*4 + z replace z with -1 x= 2 - 2*4 + -1 what is x?

OpenStudy (anonymous):

-7

OpenStudy (phi):

so that is the answer x= -7, y=4, z=-1 as a quick check, see if the 1st equation works: x + 2y – z = 2 -7 + 2*4 - (-1) = -7 + 8 + 1 = 1+1= 2 so 2 = 2 is true and it works

OpenStudy (anonymous):

Thank you so much! I feel like I actually understand how to solve these! Is there any way you can help me with one more thing? I don't know how to convert a word problem into a system of equations like this... I have an example

OpenStudy (phi):

please make it a new post

OpenStudy (anonymous):

ok